Sam and Paul have 480 marbles between them. SAm arranges them in columns with 6 in each column, and paul arranges them with 9 in each column. The total number of columns is 72. Express the above information in the form of an algebraic expression.
Answer by checkley71(8403)

X+Y=72 OR X=72-Y
6X+9Y=480
6(72-Y)+9Y=480
432-6Y+9Y=480
3Y=480-432
3Y=48
Y=48/3
Y=16 ROWS FOR PAUL'S MARBLES.
72-16=56 ROWS FOR SAM'S MARBLES.
PROOF
6*56+9*16=480
336+144=480
480=480